OpenStudy (anonymous):

NEED HELP Please! Amelia uses her credit card to buy some clothes for $678.34. She can pay off up to $235 per month. The card has an annual rate of 21.7% compounded monthly. How much total interest will she pay? $11.93 $12.27 $24.65 $29.72

jimthompson5910 (jim_thompson5910):

Month 1: Calculate Balance (by adding interest to existing balance): 678.34*(1 + 0.217/12) = 690.606648333333 = 690.61 Make payment: 690.61 - 235 = 455.61 Old Balance: 690.61 New Balance: 455.61 -------------------------------------------------------------- Month 2: Calculate Balance (by adding interest to existing balance): 455.61*(1 + 0.217/12) = 463.8489475 = 463.85 Make payment: 463.85 - 235 = 228.85 Old Balance: 463.85 New Balance: 228.85 -------------------------------------------------------------- Month 3: Calculate Balance (by adding interest to existing balance): 228.85*(1 + 0.217/12) = 232.988370833333 = 232.99 Make payment: 232.99 - 232.99 = 0 Old Balance: 232.99 New Balance: 0 -------------------------------------------------------------- Payments Made: 235, 235, 232.99 Number of Payments Made: 3 Total Cost: 235 + 235 + 232.99 = 702.99 Total Cost: $702.99 Total Interest: 702.99 - 678.34 = 24.65 Total Interest: $24.65

OpenStudy (anonymous):

Thanks! I was confused on the equation i needed to use. This helped a lot!

jimthompson5910 (jim_thompson5910):

It's not really an equation you use. It's more of the thought process you go through to get the answer